[t-interval] A random sample of size 18 is drawn from a population that is normally distributed. The sample mean is 58.5, and the sample standard deviation is found to be 11.5. Determine a 95% confidence interval about population mean.

A. [54.04,62.96]

B. [52.78,64.22]

C. [53.78,63.22]

D. [53.18,63.81]
Answer #1

Sample size = n = 18

Sample mean = = 58.5

Standard deviation = s = 11.5

We have to construct 95% confidence interval.

Formula is

Here E is a margin of error.

Degrees of freedom = n - 1 = 18 - 1 = 17

Level of significance = 0.05

tc = 2.110    ( Using t table)

So confidence interval is ( 58.5 - 5.7188 , 58.5 + 5.7188) = > ( 52.78 , 64.22)

B. [52.78,64.22]